a coin is tossed three times. write the sample space for all the outcomes.calculate the probability that the coin lands on heads at least two of those three times. show work, and write decimal to three decimal places.
Answer #1

where H=head and T=tail

Sample space S=\left \{ (H,H,H) ,(H,H,T),(H,T,H),(H,T,T),(T,H,H) ,(T,H,T),(T,T,H),(T,T,T) \right \}

The probability that the coin lands on heads at least two of those three times is \frac{4}{8}=0.5

( i.e, favorable outcomes of above event are \left \{ (H,H,H) ,(H,H,T),(H,T,H),(T,H,H) \right \} )